OpenStudy (anonymous):

What is the d value of the following sequence? 0,.5,1,1.5,2...

myininaya (myininaya):

Do you mean common difference?

myininaya (myininaya):

Just take a term and its previous term and subtract.

myininaya (myininaya):

that is all of these should be the same if the sequence has a common difference (and is aka arithmetic sequence) .5-0=? 1-.5=? 1.5-1=? 2-1.5=?

OpenStudy (anonymous):

I don't know. ;-; The question asks for the d value. I think that might be the common difference but I'm not sure.

myininaya (myininaya):

Well assuming it is the common difference, do you have anymore questions?

OpenStudy (anonymous):

Nope. ^-^ Thank you!

myininaya (myininaya):

Other than that I don't have any other guess what d value could mean.

myininaya (myininaya):

Before I leave you do you want to tell me what you think the common difference is?

myininaya (myininaya):

Any of the above subtractions thingys will give you the common difference since this is an arithmetic sequence and I know this is an arithmetic sequence because all those subtraction thingys above are the same number

OpenStudy (anonymous):

.5 It makes sense, because it is one of the answer options and it makes sense. xD

myininaya (myininaya):

yep yep :)

myininaya (myininaya):

you did awesome! :)

OpenStudy (anonymous):

Thank you! ^-^

myininaya (myininaya):

np